### v2

# Burstiness and interpersonal foraging between human infants and caregivers in the vocal domain

Abstract
Vocal responses from caregivers are believed to promote more frequent and more advanced infant vocalizations. However, studies that examine this relationship typically do not account for the fact that infant and adult vocalizations are distributed in hierarchical clusters over the course of the day. These bursts and lulls create a challenge for accurately detecting the effects of adult input at immediate turn-by-turn timescales within real-world behavior, as adult responses tend to happen during already occurring bursts of infant vocalizations. Analyzing daylong audio recordings of real-world vocal communication between human infants (ages 3, 6, 9, and 18 months) and their adult caregivers, we first show that both infant and caregiver vocalization events are clustered in time, as evidenced by positive correlations between successive inter-event intervals (IEIs). We propose an approach informed by flight time analyses in foraging studies to assess whether the timing of a vocal agent’s next vocalization is modified by inputs from another vocal agent, controlling for the first agent’s previous IEI. For both infants and adults, receiving a social response predicts that the individual will vocalize again sooner than they would have in the absence of a response. Overall, our results are consistent with a view of infant-caregiver vocal interactions as an ‘interpersonal foraging’ process with inherent multi-scale dynamics wherein social responses are among the resources the individuals are foraging for. The analytic approaches introduced here have broad utility to study communication in other modalities, contexts, and species.
Significance statement
Many behaviors—including cognitive processes—can be understood as foraging processes, with individuals balancing resource gathering with broader exploratory search. We propose that infant-caregiver vocal interactions can be understood as ‘interpersonal foraging’, with infants and caregivers seeking each others’ vocal engagement and producing more vocal activity once engagement is obtained. Over the course of the day, both human infant and adult caregiver vocalizations occur in hierarchically clustered bursts and lulls. Taking this clustering into account, we find evidence that adult responses to infants’ speech-related vocalizations promote subsequent infant vocalization and vice versa, supporting the interpersonal vocal foraging view. This provides an ecological perspective on infant-caregiver vocal interaction, connecting this communication domain to a broad array of other search behaviors.
1 Introduction
Human infants produce large numbers of vocalizations that are not yet words but are also not reflexive responses to specific emotional states (e.g., cries and laughs). These vocalizations, also known as protophones or vocal babbling, are precursors to meaningful speech [1, 2, 3]. Sometimes infants babble when they are not engaged in social interactions while other instances of babbling occur during proto-conversations between infants and adult caregivers [4], with adult vocalizations closely following infants’ and vice versa [5, 6, 7, 8, 9, 10, 4]. These proto-conversations and the contingent adult responses contained within them are understood to be important for speech and language learning [11, 12, 13, 14, 15, 16]. Longitudinal studies have also shown that caregiver responses and infant-caregiver turn-taking are predictive of later language outcomes, suggesting an enduring influence of infant-caregiver vocal interactions on language development [17, 18, 19, 20, 21].
It has been proposed that this connection between infant-caregiver vocal interactions and infant speech-language development is due to a social feedback loop wherein contingent adult vocalizations serve as rewards that guide infants’ exploratory vocal practice. The social feedback loop theory asserts that cycles of mutually contingent actions guide infant and adult vocal behavior, reinforcing infant vocalization in general and stimulating some vocalization types more than others (e.g., speech-related sounds versus cries and fusses, or more speech-like versus less speech-like protophones) [22, 6, 23, 24, 25, 15, 26]. Interactive vocal input from caregivers may be reinforcing to infants in part because it provides rich learning opportunities and infants appear to learn the value of their own vocal productions as elicitors of this social input within the first six months after birth [26].
A related way to conceptualize infant-caregiver vocal interactions is as an ‘interpersonal foraging’ process where infants and caregivers forage in the vocal domain, with sounds that elicit vocal engagement from each other serving as a possible foraged-for resource [27]. This perspective is similar to the social feedback loop theory in that social input is hypothesized to modify infant and adult vocal behavior. The foraging framing is supported by the hierarchical clustering observed in temporal patterns of both infant and caregiver vocalizations over the course of a day. Infant vocalizations and the vocal environment provided by caregiver vocalizations tend to occur in bursts of vocal activity—characterized by vocalizations in quick succession—followed by periods of silence. These bursts and lulls in turn are hierarchically clustered in time, spanning multiple timescales, and patterns of clustering are coordinated between infants and adults [28, 29]. Such hierarchical clustering is common in natural systems, and pertinently, in foraging processes. Black-browed and winged albatrosses, spider monkeys, and marine predators such as sharks have all been shown to exhibit hierarchical clustering patterns while foraging, in terms of distances (‘steps’ or ‘flights’) between stops during resource gathering, the time taken to traverse these steps (‘flight times’), or both [30, 31, 32, 33, 34]. Adult humans’ semantic memory search—in the context of a memory retrieval task that required listing as many different types of animals as possible—also showed hierarchical clustering in steps in semantic space and corresponding inter-event intervals (i.e., time between successive retrievals) [35, 36]. Further, humans’ spatial foraging patterns have been found to be correlated with their tendency toward divergent versus convergent cognitive styles [37].
Foraging across these domains has been modeled as area-restricted search (ARS). ARS is a search strategy in which an agent adaptively switches between intensive local search consisting of short spatial steps and more exploratory global search comprised of longer steps. Intensive search is utilized in response to resource encounters or expectations thereof, to effectuate exploitation in a resource-rich region while longer, exploratory steps are taken when resource encounters decline, to facilitate the discovery of resource-rich regions. Short steps correspond to short flight times—or analogously, inter-event intervals (IEIs), in the context of discrete, event-based processes—and longer steps correspond to longer flight times/IEIs [36, 38]. ARS has been shown to be an effective foraging strategy when agents are able to respond to changing environmental cues and retain memory of past events [39, 40, 41]. While human infant behavior has been explicitly characterized as foraging primarily in studies of their visual foraging via eye movements [42, 43], the idea that exploration plays a dominant role in infant behavior and learning is widely held (e.g., 44, 43). Adapting this view for the infant-caregiver vocal interaction context, IEIs for vocalization events and corresponding inter-vocalization steps (in appropriate spatial representations of the vocal domain) are expected to be shorter during times of successful interpersonal vocal engagement. Accordingly, the back-and-forth turn-taking often observed between infants and caregivers [17, 45, 46, 47, 48, 20, 21] can be viewed as periods when infants and adults are capitalizing on finding themselves in contexts of high social engagement with each other.
The idea that caregiver responses serve as motivating rewards for infant vocal babbling is also supported by computational models incorporating reinforcement-driven learning, where reinforcement by a human or simulated listener resulted in the emergence of speech-like vocal characteristics [49, 50]. In addition, experimental studies in which caregiver behavior was manipulated during short (less than 30 minutes) sessions have found that stimulating infants through contingent adult vocalizations increases the rate of infant vocalizations produced following the experimental manipulation [51, 52, 53, 5]. Moreover, selective reinforcement of infant utterances containing only vowels or containing both vowels and consonants has been found to result in a selective increase of the reinforced infant vocal type [54]. However, experimental approaches have limitations in understanding the posited bi-directional nature of infant-caregiver vocal interactions since it is not straightforward to conduct experiments to study the effects of infant responses on adult vocal behavior—it is impossible to give human infants real-time instructions on when to produce vocal responses to an adult. In addition, these experiments are limited in terms of event-level insights—information on the effect of a single instance of caregiver input on the infant’s next vocal action—and only provide session-level (e.g., over a period of several minutes) evidence that contingent responses causally increase subsequent vocalization frequency.
Figure 1: Analysis procedure. (A) shows a hypothetical sample time series of infant speech-related (ChSp; blue) and adult (Ad; black) vocalizations. ChSp inter-vocalization intervals (IEIs) are computed as the difference between the onset time of the ($i$+1)th ChSp vocalization and the offset time of the $i$th ChSp vocalization; adult IEIs are computed similarly. The response window ($T_{\rm resp}$) is indicated in purple and is used to determine whether an IEI is associated with a response or not. Child speech-related (ChSp) IEIs are excluded from response analyses (dashed gray brackets) if the IEI is less than or equal to $T_{\rm resp}$, or if the offset of the first ChSp in the sequence is followed by the onset of a child non-speech related vocalization (ChNsp; not shown) within time $T_{\rm resp}$. Otherwise, a ChSp IEI is labeled as associated with a response (shortened as ’Resp’; green brackets) if the offset of the first ChSp vocalization in the sequence is followed by an adult vocalization onset within time $T_{\rm resp}$. A ChSp IEI is labeled as not associated with a response (shortened as ‘NoResp’; red brackets) if there is no adult vocalization onset within time $T_{\rm resp}$ from the offset of the first ChSp vocalization in the sequence. Adult (Ad) IEIs (not labeled in panel A) and their associations with ChSp responses are computed similarly. (B) shows the time series of ChSp and Ad IEIs corresponding to the vocalization series in A. IEIs associated with a response are in green, IEIs not associated with a response are in red, and IEIs excluded from response analyses are in gray stripes. The arrow from ChSp IEI1 to IEI2 represents the correlation between successive IEIs. The set of equations shown summarizes the analysis steps that go into estimating this correlation and computing residual IEIs from the correlation analysis. Visualizations of these steps are provided in Fig. 2. Colors used in this figure and subsequent figures in the main text have been tested for color accessibility.
At the event level, prior research has indeed found that infant vocalizations tend to occur at higher rates than would be expected by chance following an adult vocal response and vice versa [45, 55, 10, 27]. However, these findings are confounded by the tendency for infant and adult vocalizations to occur in coordinated temporal clusters, i.e., series of vocalization events separated by relatively short IEIs. Specifically, the lack of statistical controls for such clustering means that despite the temporal proximity of infant and adult vocalizations, one or both agents may not actually be responsive to the other. For example, adult responses occurring during an ongoing infant vocalization burst can result in the misattribution of shortened infant IEIs to adult social input, even with no actual modification of infant vocal behavior as a result of adult vocal engagement. Indeed, a previous study found a strong and reliable tendency for adult vocalizations to be followed by infant cries—a finding that could be due to a tendency for cries to occur in bouts and for adults to vocalize during those cry bouts, as opposed to adult vocalizations causing infants to cry [10]. Confounds arising from inherent behavioral clustering pose unique challenges for studying the effects of resource acquisition on interpersonal foraging as conceptualized here: foraged-for resources are time-dependent and generated by the agents through their foraging behaviors, thereby making resources more likely to occur during bursts of activity.
In this study, we propose a method to address this confound and more reliably test for the effect of adult vocal input on subsequent infant vocal behavior and vice versa, by controlling for the temporal clustering of infant and adult vocalizations. We first calculate infant and adult caregiver vocal IEIs, distinguishing between IEIs based on whether the first event received a vocal response or not (see Fig. 1 and 27). We then assess the extent of temporal clustering in infants’ and adults’ spontaneous vocal productions, as evidenced by positive correlations between consecutive IEIs (Fig. 2A). Finally, we use IEI residuals from the correlation analysis to test whether adult responses predict shorter infant IEIs after controlling for infant vocal clustering (Fig. 2B) and whether infant responses predict shorter adult IEIs after controlling for adult vocal clustering. Vocal responses being predictive of shorter IEIs after controlling for correlations between consecutive IEIs would be consistent with the idea that adult vocal input following an infant vocalization serves as a reward, causing the infant to vocalize sooner than they otherwise would, and vice versa for the effect of infant vocal input on adult vocalizations. We used data from 200 daylong (10+ hour) home audio recordings collected longitudinally from infant-worn recorders when infants were 3, 6, 9, and 18 months old (totaling over 2,400 hours of audio), enabling us to test for age effects for all measures mentioned above (see Methods and Supplementary Sections S2 and S3 for details of data collection and sample sizes). Analyses were performed at the daylong and five-minute levels using automated sound segmentation and labeling and at the five-minute level using human annotations.
Figure 2: Analysis procedure. (A) is a scatter plot showing the positive correlation between successive ChSp IEIs for infants at 18 months. The infant age represented here was chosen randomly. To facilitate visualization, the plot displays 2000 randomly chosen points from the full set of daylong LENA infant IEIs at 18 months. IEIs were log-transformed and then normalized with respect to the full 18 month daylong infant IEI dataset (indicated by the $f_{z}$ function) before plotting as well as before computing residuals. The dashed line represents $\overline{\rm IEI}$, obtained by regressing current IEIs over previous IEIs within the 18 month daylong infant data (see also Fig. 1B). Note that association with response is not part of this analysis step. (B) shows the distributions of Resp and NoResp residuals ($R$) for the full daylong LENA 18 month infant IEI dataset represented in A for $T_{\rm resp}=$ 5 s. The Resp and NoResp distributions have each been normalized separately for visualization purposes. For more details on the analysis pipeline, see Methods and Supplementary Section S6.3.
2 Results
2.1 Successive IEIs are positively correlated
Figure 3: Positive correlations between previous IEI and current IEI. (A) Standardized regression coefficients ($\beta$; y-axis) from linear mixed effects analyses testing correlations between current IEI and previous IEI for infant speech-related (ChSp; blue) and adult (Ad; black) vocalizers are shown as a function of infant age in months (x-axis), for daylong vocalization data automatically labeled by LENA. 99.9% confidence intervals are shown as shaded areas and significant $\beta$ values (at $p<0.001$) are indicated as solid circles. Analyses were performed separately for each age group (3, 6, 9, and 18 months). Infant ID was treated as a random effect in all tests. (B) Standardized regression coefficients ($\beta$; y-axis) from linear mixed effects analyses testing for correlations between current IEI and previous IEI for infant speech-related (ChSp) and adult (Ad) vocalizations are shown for validation datasets: human listener-labeled 5 minute sections with all adult vocalizations included (H: All Adult; dark cyan), human listener-labeled 5 minute sections with only infant-directed adult vocalizations included (H: Child-directed Adult; green), and corresponding 5 minute sections as labeled by LENA (L: 5 min; maroon). 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p<0.001$) are indicated as solid circles. Due to limited sample size (see Supplementary Section S3 for details on sample sizes), validation data were not separated into age-blocks and instead, infant age (in months), infant ID, and infant ID-infant age interaction were treated as random effects. Note that the $\beta$ values for ChSp vocalizations from H: Child-directed Adult and H: All Adult datasets are the same, since the child vocalizations included are the same for both datasets.
For daylong LENA-segmented data, previous IEIs and current IEIs were positively correlated for both infant speech-related (ChSp) and adult (Ad) vocalizations, for all infant ages (Fig. 3A): standardized regression coefficients ($\beta$s) ranged from 0.23 to 0.30 with a mean of 0.26 for infants, and from 0.29 to 0.33 with a mean of 0.31 for adults, with $p<$ 0.001 in all cases (see Supplementary Table S7). Mixed effects regressions predicting recording-level $\beta$s—correlations between current IEI and previous IEI at the recording level—as a function of infant age found a significant negative effect with infant age for infants ($\beta=$0.01, $p<$ 0.001; Supplementary Table S9) but not for adults ($\beta=$0.003, $p=$ 0.31; Supplementary Table S9). Infant age2 effects were non-significant at the $p<0.001$ level for both infants and adults.
Fig. 3B shows the correlations between previous IEIs and current IEIs for the 5-minute excerpts segmented and labeled by human listeners, as well as for those same 5-minute excerpts’ LENA-labeled data. For the data labeled by human listeners, there are statistically significant positive correlations between previous IEIs and current IEIs for infants ($\beta=$ 0.19, $p<$ 0.001; Supplementary Table S8) and for adults, whether all adult vocalizations ($\beta=$ 0.17, $p<$ 0.001) or only infant-directed adult vocalizations ($\beta=$ 0.12, $p<$ 0.001) are considered. These results indicate that the positive correlation between successive IEIs is not an artifact of the idiosyncrasies of the LENA labeling system and is robust to labeling method.
On the other hand, the analysis of the human listener-labeled vs. LENA-labeled 5-minute excerpts suggests that the stronger current IEI $\sim$ previous IEI correlation for adults vs. infants for the daylong LENA data (Fig. 3A) is not robust to labeling method. As Fig. 3B illustrates, the correlation between current IEIs and previous IEIs is stronger for adults than infants for the LENA-labeled 5-minute excerpts while the human listener-labeled 5-minute excerpts show either no significant difference between infants and adults (when all adult vocalizations are included) or the opposite pattern (when only infant-directed adult vocalizations are included).
Age trends for infant and adult current IEI $\sim$ previous IEI correlations were not significant for the 5-minute human-labeled data (both when all adult vocalizations are considered and when only infant-directed adult vocalizations are considered) or the LENA 5-minute data (Supplementary Table S9), underscoring the lack of statistical power to resolve age trends in the validation datasets due to limited sample size (for details on sample sizes, see Supplementary Section S3).
2.2 Social inputs are associated with shorter IEIs, controlling for previous IEI
Measures of how adult caregiver responses affect subsequent infant IEIs and how infant responses affect subsequent caregiver IEIs can be expected to be confounded by the robust positive correlation between successive IEIs demonstrated in the previous section. We thus ran analyses testing whether receiving a response results in the following IEI being shorter than expected, after controlling for this positive correlation. Linear models testing for differences in IEIs associated with the receipt of a response (Resp IEIs) versus absence of a response (NoResp IEIs) were run on the residuals after predicting the current IEI based on the previous IEI. Figure 4 shows the standardized regression coefficients ($\beta$s) for these Resp vs. NoResp IEI comparisons, for response window ($T_{\rm{resp}}$) values ranging from 0.5 to 10 s. $T_{\rm{resp}}$ sets the window of time to determine response receipt as well as the threshold for the minimum IEI included in response analysis (see Methods and Fig. 1 for details). Separate linear models were run to test for each age group for LENA daylong data while validation data were not separated into age groups due to low sample size. We also ran mixed-effects regressions testing whether recording-level response effects were predicted by infant age and age2, controlling for infant ID as a random effect. However, all age effects were non-significant at $p<0.001$.
Figure 4: Response effects on IEIs after controlling for previous IEI (A) Standardized regression coefficients ($\beta$; y-axis) from linear models testing the effect of receipt of an adult (Ad) response on the length of infant speech-related (ChSp) IEIs—after controlling for the positive correlations between previous and current IEIs (see Methods and Fig. 3A for details on the current IEI $\sim$ previous IEI control)—are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis), for LENA daylong data. 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p<0.001$) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Statistical analyses were performed separately for each infant age group: 3, 6, 9, and 18 months (indicated by line and circle color; see legend). IEIs associated with responses (Resp IEIs) were significantly shorter for all age groups for $T_{\rm resp}\geq 3$ s. (B) Standardized regression coefficients ($\beta$; y-axis scale shared with panel A) from linear models analogous to A are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis) for the three different 5-minute validation datasets (see legend): human listener-labeled 5 minute sections with all adult vocalizations included (H: All Adult; dark cyan), human listener-labeled 5 minute sections with only infant-directed adult vocalizations included (H: Child-directed Adult; green), and corresponding 5 minute sections as labeled by LENA (L: 5 min; maroon). 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p$ $<$ 0.001) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Validation data were not separated into age blocks. Unlike in panel A, $\beta$s for $T_{\rm{resp}}\geq 3$ s are not significant at our $p<0.001$ threshold except for the human-labeled validation dataset in which only infant-directed Ad vocalizations are included (see Supplementary Figs. S30 and S31 for significant $\beta$ values at less stringent significance thresholds). However, they do trend negative for all three validation datasets for $T_{\rm{resp}}\geq 3$ s. (C) $\beta$ values (y-axis) from a similar model as in A, testing the effect of infant response on adult IEI length are shown for LENA daylong data. Resp IEIs are significantly shorter for all age groups for $T_{\rm resp}\geq 2$ s. (D) $\beta$ values (y-axis shared with panel C) from a similar model as in B testing the effect of infant response on adult IEI length are shown for validation datasets. For human-labeled vocalizations, when only infant-directed adult vocalizations are included, there is a significant negative effect of infant response on adult IEI for 1 s $\leq T_{\rm{resp}}\leq 5$ s.
2.2.1 Effect of adult responses on infant IEIs
For $T_{\rm{resp}}\geq 2$ s, infant IEIs were significantly shorter ($p<0.001$) after receiving an adult response for LENA daylong data, as evidenced by the negative response $\beta$s in Fig. 4A. This pattern of negative $\beta$s holds for all age groups for full day LENA-segmented data, except for 6 month-old infants at $T_{\rm resp}=2$ s, in which case the trend was consistent but did not reach significance at $p<0.001$ (see Supplementary Fig. S30 for significant $\beta$ values at less stringent significance thresholds). For the representative case of $T_{\rm resp}=5$ s (see Fig. 2B), $\beta$s ranged from -0.09 to -0.15 across all age groups (see Supplementary Table S10 as well as the OSF repository associated with this study).
For validation datasets using human-labeled 5-minute excerpts (Fig. 4B), the negative $\beta$ was significant only for the human-labeled dataset with only infant-directed adult vocalizations included at $T_{\rm resp}=3$ s. All other $\beta$s trended in a direction consistent with the negative $\beta$s found for the daylong LENA-segmented data for $T_{\rm{resp}}\geq 2$ s but were not statistically significant at the $p<0.001$ threshold (see Supplementary Fig. S31 for significant $\beta$ values at less stringent significance thresholds and Supplementary Table S12 for $\beta$ values for the representative case of $T_{\rm resp}=5$ s). The corresponding LENA-labeled 5-minute portions also had non-significant negative response $\beta$s for $T_{\rm{resp}}\geq 2$ s, suggesting that the non-significance may primarily be an issue of insufficient statistical power with these much smaller datasets.
For $T_{\rm resp}<2$ s, IEIs were longer after an adult response for LENA daylong data, as evidenced by positive $\beta$ values for all age groups. This trend was significant at $p<0.001$ at $T_{\rm resp}=0.5$ s for infants at 6, 9, and 18 months, and at $T_{\rm resp}=1$ s for 6 month old infants. The tendency for adult vocalizations to be longer than most other sound types (see Supplementary Table S2 and Supplementary Fig. S2) may contribute to positive $\beta$ values at $T_{\rm resp}<2$ s, which correspond to cases where the shortest IEIs in the data are included in response analyses (see also Supplementary Fig. S27). IEIs associated with a response necessarily have an intervening adult vocalization. IEIs not associated with a response, on the other hand, can only have silence or sound types other than infant (speech and non-speech related) or adult—which may generally be shorter than adult vocalizations—intervening. At low response window thresholds, short IEIs for which the limiting factor on IEI length is the intervening sound type are more likely to be included in the analysis. As a result, for small $T_{\rm resp}$ values, Resp IEIs may tend to be longer than NoResp IEIs purely as a function of the duration of the intervening adult response vocalization. This pattern of positive $\beta$s for $T_{\rm resp}<2$ s also holds for the validation datasets, with significant $\beta$s at $T_{\rm resp}=0.5$ s for all three validation datasets and at $T_{\rm resp}=1$ s for the human-labeled 5-minute excerpts with all adult vocalizations included. Thus, as long as the response window parameter, $T_{\rm{resp}}$, is sufficiently long, results are consistent with a tendency for adult vocal responses to promote infant vocalization, causing infants to vocalize again sooner than would be expected based on the infant’s previous inter-vocalization interval.
2.2.2 Effect of infant response on adult IEIs
Adult IEIs were shorter following an infant response for $T_{\rm{resp}}\geq 2$ s. This effect is significant for all infant age groups for the full-day LENA-labeled data (Fig. 4C). $\beta$s ranged from -0.11 to -0.17 across all infant age groups for the representative case of $T_{\rm resp}=5$ s (see Supplementary Table S11).
For the human-labeled 5-minute datasets (Fig. 4D), the response effect is significant for 1 s $\leq T_{\rm{resp}}\leq 5$ s when only infant-directed adult vocalizations are included in the analysis, with the significant effect sizes being stronger than the effect sizes for the human-labeled 5-minute datasets with all adult vocalizations included, indicating that infant responses affect infant-directed adult vocalizations more than non-infant-directed adult vocalizations. For human-labeled 5-minute portions with all adult vocalizations included and for LENA labels for those same 5-minute portions, response effects are significant for a more limited range of $T_{\rm{resp}}$ values.
For very short $T_{\rm resp}$ values ($<2$ s), response effects for the daylong LENA-labeled data are non-monotonic as a function of $T_{\rm resp}$: $\beta$s for $T_{\rm resp}=0.5$ s are negative and statistically significant at $p<0.001$ for all age groups while $\beta$s for $T_{\rm resp}=1$ s are greater than those at $T_{\rm resp}=0.5$s but only significant at 18 months (Fig. 4C). This non-monotonicity is not mirrored in the human-labeled validation data. Therefore, it is possible that the pattern of increase followed by decrease observed for LENA daylong response $\beta$s at very short response windows is an artifact resulting from a combination of LENA’s minimum duration thresholds for different sound types and differences in typical vocalization durations.
LENA child vocalizations have a 600 ms minimum and modal duration (see Supplementary Figs. S2, S13), which is shorter than the 800–1000 ms minimum durations for all other sound type labels as reported in the LENA documentation [2] (our data indicate that most sound types—including adult vocalizations—have durations lower than these published minima; see also Supplementary Table S2 and Fig. S2). Thus, for adult IEIs at $T_{\rm{resp}}=0.5$ s, it may be more likely that only one sound type intervenes, since IEIs as short as $T_{\rm{resp}}$ are included in the analyses. For Resp IEIs, this sound type must, by definition, be the target child, which can be expected to be relatively short, resulting in short Resp IEIs. However, when $T_{\rm{resp}}=1$ s, Resp IEIs associated with a relatively short infant response may have a second intervening sound type—since IEIs must be greater than $T_{\rm{resp}}$ to be included in the analysis—and this could lead to Resp IEIs being longer. For the validation data, response effect patterns at low $T_{\rm resp}$ values appear to be similar to results for infant IEIs, supporting the possibility that response effects at low $T_{\rm resp}$ values for the daylong data may be primarily driven by LENA’s minimum duration thresholds for different sound types.
For the research questions that are the focus of this study, the results are most stable and interpretable for $T_{\rm resp}\geq 2$ s. When this is the case, for adult IEIs, similar to infant IEIs, results are consistent with a tendency for infant vocal responses to promote adult vocalization, causing adults to vocalize again sooner than they otherwise would have.
3 Discussion
The modification of foraging trajectories in response to resource acquisition is a key feature of adaptive foraging processes such as area-restricted search (ARS). The adaptive hallmark of ARS is a switch from exploratory search consisting of longer steps and inter-event intervals to exploitative search involving shorter steps and inter-event intervals in response to resource acquisition. This trade-off between exploitation and exploration is observed in adaptive foraging processes in physical space as well as in cognitive domains, where search processes aren’t anchored to an explicit physical space [57, 31, 58, 59]. In the context of framing infant-caregiver vocal interactions as an interpersonal foraging process, we posit that the resources being sought for include vocal engagement from each other [27]. However, prior research looking at infant vocal behavior following adult vocal input has not been able to resolve whether adult social responses causally increase the likelihood of subsequent infant vocalization or whether adult responses are merely embedded within clusters of infant vocalizations that occur as part of the bursts and lulls in infants’ spontaneous vocal production, without infants actively modifying their behavior as a result of receiving adult input [10].
To address this confound, we analyzed infant and adult IEIs, informed by the analysis of step sizes and flight times in foraging studies. We primarily used longitudinal data consisting of infant-centered daylong audio recordings automatically labeled by the LENA system. A subset of 5-minute excerpts from the daylong data labeled by human listeners and LENA-labeled data for the same 5-minute excerpts served as validation datasets. We first established that successive infant and adult IEIs are indeed positively correlated, for both the full-day automatically labeled dataset and the validation datasets. We also found evidence that successive infant IEIs were less correlated as infant age increased for the LENA daylong data, pointing towards increased baseline exploratory vocal practice by infants as they get older (see also Supplementary Fig. S9B). This trend, however, was not statistically significant for the human-labeled validation data and likely requires more validation data for verification.
Correlation between successive IEIs as demonstrated in this study serves as a metric for assessing clustering that complements other measures of temporal clustering in infant vocalizations [28, 60] as well as other behavioral domains such as language production (e.g., 61), infant play behavior (e.g., 62), and animals’ foraging in space (e.g., 31). This underscores the broad potential utility of correlations between consecutive IEIs as a measure of event-level clustering in assessing the temporal structure of event time series. A benefit of this successive IEI correlation metric in studying interactive behavioral time series (such as infant-caregiver vocal interactions) is that it provides a method to control for temporal clustering when looking for immediate effects of social engagement.
We defined social engagement as any vocalization onset from the communication partner occurring within a specified response window ($T_{\rm{resp}}$, seconds) following the offset of the focal individual’s vocalization. We considered $T_{\rm resp}$ values ranging from 0.5 s to 10 s which allowed us to test for response window values for which a reliable social engagement effect could be detected. Finally, we tested whether residual IEIs after predicting current IEI based on prior IEI were smaller if the first vocalization comprising the IEI was followed by interpersonal vocal engagement. These analyses revealed that infants vocalized sooner following an adult response across all age groups in the automatically labeled data for $T_{\rm{resp}}\geq 3$ s. We found similar patterns for adults: adult IEIs were significantly shorter following an infant response across infant ages for the LENA daylong data for $T_{\rm{resp}}\geq 2$ s. Results from analyses of the validation datasets supported these findings, although not all validation results reached statistical significance.
This pattern of shorter IEIs following social input breaks down when the $T_{\rm{resp}}$ parameter is very small; in the present study, this was about 1 s or lower. This may be due to the fact that very short response windows allow for the inclusion of very short IEIs—the length of which may be dictated primarily by the duration of the intervening sound type—in the analysis. Using larger $T_{\rm{resp}}$ values also provides increased flexibility in capturing responses that are delayed due to competing demands on caregivers’ attention, or that are effortful and take time to activate and execute, especially for infants at early stages of speech production [46, 14]. Further, since our method does not distinguish between effects that are specific to contingent responses versus more general stimulation as a result of social input, larger $T_{\rm{resp}}$ values will include more of the latter type of events, resulting in increased sample size for IEIs associated with a response and consequently, statistical power. On the other hand, the longer the response window, the more IEIs are excluded from analysis (see Supplementary Section S6.3). Performing a sweep across a range of $T_{\rm{resp}}$ values as we did in the present study allows for empirically choosing suitable values of $T_{\rm{resp}}$ to achieve balance across these considerations.
Taken together, our findings support the framing of infant-caregiver vocal communication as an interpersonal foraging process, indicating that social engagement from the other agent promotes subsequent vocal activity by both infants and adults in real-world contexts experienced over the course of a day. Within the interpersonal foraging framework, our results are readily analogous with prior work on area-restricted search, showing that the distance traveled—and time elapsed—between foraging stops tends to be shorter when the foraging agent is retrieving resources at a high rate and longer when resource obtainment slows [36, 38].
The pattern of responses being associated with shorter IEIs could reflect a process of learning (e.g., learning to expect responses) or a general tendency for vocal input from others to stimulate infant (and adult) vocal activity. This has frequently been noted as a limitation of experimental studies manipulating adult responses to infant vocalizations [51, 52, 63, 5]. Indeed, at least one experiment has found that a general vocal elicitation effect of adult input influenced infant volubility more than temporal contingency of the adult input immediately following infants’ vocalizations [5]. This ambiguity also applies in the case of the effects of infant responses on adult IEIs. The application of our analysis approach to experimental or simulation data could help assess and refine the approach with respect to this distinction.
Another limitation of our approach is that it only considers the most recent IEI when predicting current IEI. In the future, incorporating more vocalization history may further enhance the prediction of current IEI based on prior behavior, which may in turn amplify the true signal when testing for the effects of receiving social input. The current approach also does not address the possibility of vocal extinction bursts—a phenomenon observed in the still face experimental paradigm where infants exhibit a sequence of increased then decreased vocal activity following cessation of engagement by an adult [64, 65, 26]. How reward extinction manifests in infants’ real-world experiences and how to identify instances of extinction bursts in real-world interactions remain open questions.
Finally, interpersonal foraging can be compared to other types of social foraging, such as when there are cooperative or competitive interactions between foragers [66, 67, 68]. In both cooperative and competitive foraging, the occurrence of the foraged-for resource is independent of the foraging agents, even as other agents’ behavior inform foraging strategy. In contrast, resources in interpersonal foraging are generated by the agents’ foraging behaviors. This is similar to some cases of foraging for mates in that the agents are searching for each other and foraging interactions between agents are not antagonistic. The diversity of social foraging processes notwithstanding, dynamic interactions between foraging agents is a common thread shared by social foraging and interpersonal foraging as conceptualized here. As such, it could be fruitful to adapt computational models of social foraging to interpersonal foraging and assess implications for optimal foraging strategies.
In the future, this approach could be used to compare child-caregiver vocal interactions across diverse cultural and socioeconomic contexts [69]. Moreover, the burstiness of communicative signals and the importance of social vocal engagement in infant vocal development are shared by a number of animal communication systems [70, 71, 72, 73, 74]. As such, the use of correlations between successive IEIs to control for temporal structure in behavioral event series as well as the interpersonal foraging framework proposed in this study could be applied to understand interactive behaviors in other animal species [75], such as marmosets, which also exhibit infant-caregiver vocal turn-taking [76], as well as in other communication modalities [16, 77], such as gestures in humans [78, 79, 80] and great apes [74].
In summary, the current study provides robust evidence for facilitatory effects of social engagement on both infant and adult vocalizations in naturalistic, ecologically relevant contexts, after controlling for the inherent burstiness of vocal behavior. These findings support the view that infant vocalization is an interpersonal foraging process.
4 Methods
4.1 Data collection
Data used in this study consists of daylong (10+ hour) naturalistic recordings (recorded on a typical day in the infant’s home environment) collected longitudinally from 54 infants aged 3, 6, 9, and 18 months using the LENATM Pro System. 45 infants had recordings at all four ages (see Supplementary Section S2 for further details). Infants resided in the San Joaquin Valley region of California and for the most part were exposed primarily to English and/or Spanish, although other languages were also spoken in some of the infants’ homes. Socioeconomic status (SES) of caregivers varied: seven households reported an annual income of less than \$30,000, seven reported between \$30,000-60,000, six reported between \$60,000-90,000, and 9 reported over \$90,000. Highest level of education within the families also varied: 2 completed some high school, 3 completed high school or received a GED, 10 completed some college, 3 completed an associate’s degree, 18 completed a bachelor’s degree, 11 completed a master’s degree, and 7 completed a doctoral degree. Infants’ reported ethnicity/race were as follows: 17 White, 13 Hispanic/Latino (race not specified), 13 Hispanic/Latino - White, 2 Hispanic/Latino - Native American, 2 Black/African American, 1 Asian, 4 Asian and White, and 2 Unknown. Caregivers were instructed to turn recorders on and put them on the infants by 8am and to finish recording no earlier than 7pm. Caregivers were allowed to pause the recorder during the day as needed for privacy purposes and multiple pauses were allowed. However, the total pause time was not to exceed 1 hour for a recording day. Caregivers were also able to request sections of the recording be deleted for privacy purposes. Monetary compensation was provided to participants based on the completion of each recording and associated questionnaires.
This is an expanded version of the dataset used in a previous study [27], which can be referred to for further details regarding the data collection methods.
4.2 Identification of infant and adult vocalizations
Audio recordings were processed by the LENA Pro software, which provided vocalization onset and offset times as well as segment labels. Our analyses are primarily focused on segments labeled ‘CHN’ (defined as clear vocalizations likely produced by the infant wearing the recorder) containing at least one ‘Utt’ (a speech-related vocalization, as opposed to cry, laugh, or vegetative vocalizations), which we refer to as ‘ChSp’ (short for Child Speech-related); ‘FAN’ (clear adult female vocalizations); and ‘MAN’ (clear adult male vocalizations). FAN and MAN vocalizations were grouped into the ‘Ad’ (short for Adult) category for the purposes of our analyses, and collectively represent potentially as many adults as were in the infant’s vicinity over the course of the day [2]. We use the short-hand ‘ChNsp’ (short for Child Non-speech-related) to refer to CHN segments that do not contain an Utt and thus, are considered likely to be only cry or vegetative. ChNsp segments were not analyzed in this study but were sometimes used to determine which inter-event intervals (IEIs) were eligible for analysis (see below). Unless otherwise specified, all references to infant vocalizations indicate ChSp vocalizations in the context of this study. For ChSp, ChNsp, and Ad vocalizations, all instances of more than one vocalization of the same category separated by 0 s were merged to create a single vocalization (see Supplementary Section S6.2 for details). Note that FAN and MAN vocalizations separated by 0 s were merged into a single Ad vocalization under this protocol.
4.3 Defining response (Resp) and no-response (NoResp) inter-event intervals (IEIs)
Infant IEIs were defined as the time between the offset of a ChSp vocalization and the onset of the next ChSp vocalization. Adult IEIs were defined as the time between the offset of an Ad vocalization and the onset of the next Ad vocalization. Recorder pauses during a recording day and deletions of portions of recorded audio created multiple sub-recordings. We did not compute IEIs that spanned sub-recordings.
An infant (ChSp) vocalization was deemed to have received a caregiver response if the infant vocalization offset was followed by an adult vocalization onset within the response window threshold, $T_{\rm resp}$, with no intervening ChSp or ChNsp vocalizations (Fig. 1). In this case, the IEI from the offset of the ChSp vocalization that received the response to the onset of the next ChSp vocalization was considered a with-response (‘Resp’) IEI. If an infant vocalization offset was followed by the onset of an infant speech-related (ChSp) or non-speech-related (ChNsp) vocalization within time $T_{\rm resp}$ without an intervening adult vocalization, the response was coded as ‘Not Applicable’ (NA). IEIs associated with NA responses were excluded from analyses testing the effect of response reciept. Finally, infant vocalizations whose offsets were not followed by the onset of an adult vocalization or either infant vocalization types (ChSp, ChNsp) within time $T_{\rm resp}$ were considered to not have received a caregiver response, and the corresponding IEIs were considered without-response (‘NoResp’) IEIs. Resp IEIs $\leq T_{\rm resp}$ were excluded from analyses comparing Resp IEIs to NoResp IEIs to avoid the confound of Resp IEIs being shorter due to not needing to wait the full $T_{\rm resp}$ time period to determine response receipt. $T_{\rm resp}$ ranged from 0.5 s to 10 s, with the range between 1 s and 10 s spanned by 1 s increments. Estimating infant response receipt to adult vocalizations followed an analogous approach, swapping ChSp with Ad roles in the procedure described above.
4.4 Statistical tests relating IEI length to response receipt
Statistical analyses were performed separately for ages 3, 6, 9, and 18 months to assess whether response receipt affected IEIs at each age group. All IEIs were log-transformed (base 10) and then normalized prior to subsequent analysis steps. Normalization was performed with respect to the age-block subset (e.g., infant IEIs at 3 months) being analyzed and did not take response receipt into account.
To test whether receiving an adult response had an effect on infant vocalization IEIs, we used a two-step approach designed to control for the endogenous clustering of infant (or adult) vocalizations in time (Fig. 2). Separate tests were conducted for each $T_{\rm resp}$ value. First, a linear mixed effects model was built, predicting current infant IEI with the previous infant IEI as a fixed effect and infant ID as a random effect. Next, residuals from this model were used as the dependent variable in a linear model that had one binary fixed effect representing whether the residual was from a Resp or a NoResp IEI. To test the effect of infant age on how response receipt affected IEIs, the above protocol was carried out for each recording day, followed by a linear mixed effects model predicting the resultant recording day-level response effect $\beta$s, with infant age and age2 as fixed effects and infant ID as a random effect. Analogous analyses were carried out for adult IEIs and infant responses.
Because of the large number of statistical tests performed, we opted to only interpret effects with $p<0.001$ to be statistically significant.
4.5 Validation using 5-minute excerpts labeled by human listeners
To validate the results based on LENA’s automated infant and adult vocalization tagging, we used a subset of the LENA daylong data from 45 infants, in which 420 5-minute sections from 148 daylong recordings (up to three 5-minute sections per recording) were independently segmented and labeled by trained human listeners (see Supplementary Section S2 for further details). These 5-minute sections were selected on the basis of having high numbers of child vocalizations (i.e., high child volubility) as suggested by the LENA software and representing different portions of the day. Section selection involved a criterion of the sections being 30 minutes apart, although some 5-minute sections were less than 30 minutes apart due to human error (see Supplementary Section S2 for details). Research assistants used ELAN [81] and followed a protocol that asked them to label the onsets and offsets of all target infant and adult vocalizations and to identify major sub-types of each. For infant vocalizations, these sub-types were canonical (C), non-canonical non-reflexive (X), cry-reflexive (R), and laugh-reflexive (L). The C and X sub-types together form an equivalent class to the CHN Utt label by LENA and therefore, our ChSp category, while the R and L sub-types together correspond to our ChNsp category. For adult vocalizations, sub-types indicated whether the vocalization was directed to the target child (T), not to the target child (N), or addressee unknown (U). Custom R and MATLAB scripts were written to check for potential annotation errors, such as putting annotations in the wrong tier, mistyped or missing annotations, and annotations with no (or negative) duration. All identified errors that could not be corrected automatically were corrected by one of the authors (JM) who was also one of the original trained human listeners. We also identified two cases where the annotation file was linked to the incorrect audio file. Affected annotations were discarded and excluded, and 5-minute sections from the discarded files were not included in the set of 420 5-minute sections we analyzed in this study.
Unlike the LENA system’s labels, infant and adult vocalizations identified by human listeners could overlap. For consistency with the LENA system, these overlaps were excluded from further analyses (see Supplementary Section S6.1 for details). However, there were still some differences from how the LENA system treats overlaps. Since human listeners only tagged adult and infant vocalizations, only infant-adult overlaps were excluded; cases of overlaps of either adults or the target infant with other children, electronic sounds, or environmental noise were not excluded. In addition, unlike LENA labels, the human listener labels did not have prescribed duration minima (see Supplementary Sections S3.2, S3.3, and S4.2).
To test the validity of the LENA-based analyses, we considered three different sets of labels for these 5-minute high-infant-volubility sections: (1) the automatically generated LENA labels for these sections (L: 5 min), (2) the set of human-tagged infant speech-related (ChSp) and non-speech related (ChNsp) vocalizations together with the set of all human-tagged adult vocalizations (H: All Adult), and (3) the set of human-tagged ChSp and ChNsp vocalizations together with the set of human-tagged infant-directed adult vocalizations (H: Child-directed Adult). Together, these three sets of labels allowed us to identify consistencies and differences in results across labeling methods, controlling for the portions of the day that are sampled.
IEIs from the three validation datasets (L:5 min, H: All Adult, and H: Child-directed Adult) were analyzed following the same procedures employed for the full-day LENA-annotated dataset, except that data were not separated by infant age and infant age was considered as a random effect rather than a fixed effect. Infant ID and infant ID-infant age interaction were used as additional random effects. Additional overall reliability metrics for the human versus LENA labeling are provided in Supplementary Section S5.
Acknowledgments
We thank the families who participated, the many research assistants who helped with data collection and annotation, and Gina Pretzer and Jimel Mutrie for coordinating, supervising, and performing data collection and annotation. This work was supported by the National Science Foundation (grant numbers 1529127 and 1539129/1827744) and by the James S. McDonnell Foundation (https://doi.org/10.37717/220020507).
Ethics information
All data for this study was collected in accordance with relevant guidelines and regulations. Informed consent was obtained from the infant participants’ legal guardians. Data collection protocols were approved by the University of California, Merced Institutional Review Board (protocol ID: UCM15-0031) and continuing data analysis was approved by the University of California, Los Angeles Institutional Review Board (protocol IDs: IRB-23-1317 and IRB#23-001317).
Statement of Authorship
ASW, VPSR, CTK, and AG designed the research; SS coordinated, SS and LL performed, and ASW and CTK supervised data collection; JM performed, and SS, LL, and ASW supervised data annotation; VPSR and JM cleaned and pre-processed data; VPSR performed statistical analyses; VPSR and ASW wrote the paper with input from CTK, AG, LL, and SS.
Data and Code availability
All processing and analysis code is available on Github. De-identified preprocessed data used in the present paper as well as all relevant results tables are available on OSF. The raw audio data and LENA labels for participants who gave permission for us to share their data with other researchers are available within the San Joaquin Valley corpus [82] in HomeBank [83].
Supplementary Information
S1 Abbreviations for the datasets
  • L (day): LENA daylong data consisting of daylong recordings collected, automatically segmented, and labeled using the LENA™ system.
  • H (5min) and H (5min; All Ad): Validation data consisting of human listener-annotated 5 minute sections of L (day) data, with all adult vocalizations included. Note that H (5min) and H (5min; All Ad) abbreviations are used interchangeably. In the main text, this dataset is referred to as H: All Adult.
  • H (5min; T-Ad): Validation data consisting of human listener-annotated 5 minute sections of L (day) data, with only infant-directed adult vocalizations included. The ‘T’ refers to the annotation label for infant-directed adult vocalizations, specifying the adult vocalization as addressed ‘to’ the infant wearing the recorder. In the main text, this dataset is referred to as H: Child-directed Adult.
  • L (5min): LENA-labeled counterpart of the human listener-labeled validation data, consisting of LENA segment onsets, offsets, and labels for the 5-minute sections annotated by human listeners. In the main text, this dataset is referred to as L: 5 min.
S2 An overview of the data: number of files [^1] Note that the terms ‘file’ and ‘recording’ are used interchangeably in this document. and contributing infant ages
The daylong data used in this study consists of 200 daylong recordings from 54 infants. Recordings were collected at ages 3, 6, 9, and 18 months; 45 infants have recordings at all four ages.
The validation data consists of 420 5-minute sections from 148 daylong recordings from 45 infants. 30 infants have validation data at all four ages. Up to three 5-minute sections were selected for annotation from each daylong recording based on the LENA system identifying them as having high child volubility. 18 recordings in the validation dataset do not have three annotated sections. Of these 18 files, one file (infant ID 225, 3 months) has four 5-minute sections while all others have fewer than three 5-minute sections annotated. One section (infant ID 014, 3 months) chosen for annotation was 3 minutes long (instead of 5 minutes) due to human error. 5-minute sections were assigned to human annotators in three batches, as annotation began prior to the completion of data collection. Within each batch, sections were assigned in random order. Some of the assigned 5-minute sections were not included in the validation dataset because the human annotation phase of the project ended before these sections were completed. While most validation files have the 5-minute sections at least 30 minutes apart, 17 files have a pair of 5-minute sections less than 30 minutes apart due to human error.
For a breakdown of the number of recordings in the LENA daylong dataset and the validation dataset by infant age, see Table S1. Each month-level infant age group corresponds to a distribution of infant ages at the day-level. These day-level infant age distributions for LENA daylong data and validation data are shown in Fig. S1.
Infant age (months) Number of recordings: LENA daylong data Number of recordings: Validation data
3 54 43
6 51 35
9 50 37
18 45 33
Total 200 148
Table S1: Breakdown of number of recordings by infant age for LENA daylong data and validation data. The last row of the table is the total number of recordings (summed across infant age) by dataset type.
Figure S1: Distribution of infant ages (in days) in LENA daylong data and validation data. (A) shows the number of recordings at each unique infant age (in days) in the LENA daylong data. The bars are color-coded by the month-level age group (3, 6, 9, or 18 months; see legend) the recordings belong to. (B) shows the number of recordings at each unique infant age (in days) in the validation data, color coded by month-level age group. A and B share a legend.
S3 Numbers and durations of audio segments by type
S3.1 Overview
The LENA daylong data amounts to a total of 2418.10 hours of recorded audio, of which 571.12 hours (23.62% of total recorded time) are silence (SIL), while key segments of interest—Child Speech-related (ChSp), Child Non-speech-related (ChNsp), and Adult (Ad)—cumulatively amount to 329.25 hours (13.62% of total recorded time). Mean recording length is 12.09 hours, with a standard deviation of 1.78 hours and a range of 9.33–16.00 hours.
The validation data amounts to a total of 34.97 hours of recorded audio from a total of 420 annotated sections. All sections are 5 minutes long except for one which was only 3 minutes long due to human error in generating the annotation assignments. For L (5min) data, key segments make up 11.21 hours (32.60% of total recorded time) while for H (5min) data, key segments make up 11.87 hours (33.91% of total recorded time), after removing overlaps (see Methods and Section S6.1). For H (5min; T-Ad) data, key segments make up 9.25 hours (26.44% of total recorded time) after removing overlaps, reflecting the removal of adult vocalizations that are not directed towards the key infant (U and N annotations), which make up 2.61 hours (7.47% of total recorded time).
S3.2 Summary statistics: LENA daylong data
Segment label Description Mean segment duration (s)
ChSp Key child, speech-related 1.14
ChNsp Key child, non-speech-related 1.43
CHF Key child, faint 0.80
Ad Female or male adult 1.71
FAF Female adult, faint 1.19
MAF Male adult, faint 1.42
CXN Other child 1.13
CXF Other child, faint 1.01
OLN Overlap 2.28
OLF Overlap, faint 1.60
TVN Electronic 2.14
TVF Electronic, faint 1.56
NON Noise 4.31
NOF Noise, faint 1.78
SIL Silence 3.18
Table S2: Mean segment duration for LENA daylong data. Mean durations (in seconds) for all segment labels as identified by LENA™ for the LENA daylong data are shown. Key segment labels (infant speech-related, ChSp; infant non-speech-related, ChNsp; Adult, Ad) are indicated in bold. Segment labels are organized such that faint sounds of each segment label category follows the ‘near’ segment label for that category. For example, CHF is the segment label for very quiet instances of vocalizations by the child wearing the recorder. Similarly, FAF and MAF are respectively faint female and faint male adult, while Ad is combined female or male adult vocalizations that are louder from the infant’s perspective (LENA’s ‘FAN’ and ‘MAN’ categories, respectively, combined into a single adult category). ChSp, ChNsp, and Ad summary statistics are computed after merging segments with the same label separated by 0 s IEIs (see Methods for more details). For more details on the segment labels and descriptions, see LENA technical reports [1] and [2]. The numbers in this table correspond to the closed circles in Fig. S2B.
Figure S2: Summary statistics for number of segments in a recording and segment duration for LENA daylong data. (A) Mean (closed circle) and median (open square) number of segments in a recording day for all segment labels identified by LENA are shown. Error bars indicate minimum and maximum values across all recordings. Key segment labels (infant speech-related, ChSp; infant non-speech-related, ChNsp; Adult, Ad) are shown in red, silence (SIL) is shown in grey, and all other segment labels are in black. For sound segment labels, all ‘near’ (N) sounds are indicated by solid lines while all faint (F) sounds are indicated by dashed lines (see [1] and [2] for definitions of near and faint). Note that ChSp, ChNsp, and Ad are all near sounds. (B) shows the mean (closed circle) and median (open square) individual segment durations (in seconds) for all segment labels identified by LENA. A and B share a legend. For A and B, ChSp, ChNsp, and Ad summary statistics are computed after merging segments with the same label separated by 0 s IEIs (see Methods for more details). Summary statistics for all other segment labels are presented without similar merging. We make this distinction because the 0 s IEI merging is part of the processing pipeline for ChSp, ChNsp, and Ad segments with respect to the results presented in this paper. As such, we wanted to preserve that information here.
S3.3 Summary statistics: Validation data
Segment label Mean duration (s): L (5min) Mean duration (s): H (5min)
ChSp 1.42 0.78
ChNsp 1.53 1.38
Ad 1.55 1.10
Table S3: Mean segment duration for validation data. Mean durations (in seconds) for key segment labels (infant speech-related, ChSp; infant non-speech-related, ChNsp; Adult, Ad) for L (5min) data and H (5min) data are shown. Means are computed after merging segments with the same label separated by 0 s IEIs (see Methods for more details) and, for H (5min) data, after overlaps have been removed (see Methods and Section S6.1 for details).
For human listener-labeled data with only child-directed adult vocalizations included, i.e., H (5min; T-Ad), the mean duration of adult vocalizations is 1.03 s—slightly lower than the mean adult vocalization duration when all adult vocalizations are included (Table S3). Mean ChSp and ChNsp durations are the same for H (5min; T-Ad) and H (5min) data since all infant vocalization segments are the same for these two datasets (see also Fig. S3).
Figure S3: Summary statistics for number of segments in a file and segment duration for validation data. (A) Mean (closed circle) and median (open square) number of segments in a validation data file for key segment labels (ChSp, ChNsp, Ad) are shown for L (5min) data (black), H (5min) data (green), and H (5min; T-Ad) data (blue). Summary statistics from H (5min; T-Ad) data is only shown for Ad segments because all infant vocalizations—and therefore, associated summary statistics—are the same for H (5min; T-Ad) and H (5 min) datasets. Error bars indicate minimum and maximum values across all files and each file contains up to four 5-minute sections. (B) Mean (closed circle) and median (open square) individual segment durations (in seconds) for key segment labels are shown for L (5min) data (black), H (5min) data (green), and H (5min; T-Ad) data (blue). A and B share a legend. Summary statistics in A and B are computed after merging segments with the same label separated by 0 s IEI (see Methods for more details), and in the case of human listener-labeled data, after removing overlaps (see Methods and Section S6.1 for details).
S3.4 LENA daylong and validation data compared
Figure S4: Vocalization number and duration totals for LENA daylong data and validation data. (A) Total number of vocalizations (sum of the number of adult and infant—speech-related and non-speech-related—vocalizations; $n_{\rm{tot}}=n_{\rm{Ad}}+n_{\rm{ChSp}}+n_{\rm{ChNsp}}$) are shown for LENA daylong data (L (day), yellow), human listener-labeled 5-minute sections (H (5min), green), and corresponding 5-minute sections as labeled by LENA (L (5min), maroon). (B) Total duration of vocalizations (sum of the duration of adult and infant vocalizations; $d_{\rm{tot}}=d_{\rm{Ad}}+d_{\rm{ChSp}}+d_{\rm{ChNsp}}$) are shown for L (day), H (5min), and L (5min) data. For A and B, the bars for H (5min) data represent totals after overlaps have been processed. For all three datasets, totals are presented after 0 s IEIs have been merged.
For human listener-labeled data with only infant-directed adult vocalizations included (H (5min; T-Ad)), the total number of key segments (ChSp, ChNsp, and Ad) is 37377, while the total duration of key segments is 33319.553 s. See Figs. S6 and S8 for details.
Figure S5: Number of vocalizations in LENA daylong data by infant age and vocalization type, expressed as totals at the dataset-level and as box plots for totals at the recording day-level. (A) Total numbers of infant (Ch; dark blue) and adult (Ad; green) vocalizations in the L (day) dataset for each infant age (months; x-axis) are shown as a bar plot. Bars for infant speech-related (ChSp) and infant non-speech-related (ChNsp) totals are shown within the infant (Ch) bars, indicating that these are sub-categories. (B) Box plots summarizing total numbers of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations at the recording day-level are shown as a function of infant age (months; x-axis). Box plots whose notches do not overlap have different medians at the 0.05 significance level [^2] Per the documentation of MATLAB’s (R2024b) boxchart function, which was used to generate box plots, “the significance level is based on a normal distribution assumption, but the median comparison is reasonably robust for other distributions”.. Number of vocalizations are reported after vocalizations with the same label are merged when separated by 0 s IEIs. Note that total vocalization numbers shown in A do not account for the fact that different ages have different numbers of daylong recordings (see Table S1). As such, B provides a more accurate representation of how the total number of each vocalization type in a daylong recording changes with infant age. However, data plotted in B have not been normalized to account for differences in total number of vocalizations at the recording-level due to variations in recording duration across daylong recordings (see Section S3.1).
Figure S6: Number of vocalizations in the validation datasets by vocalization type, expressed as totals at the dataset-level and as box plots for totals at the recording-level. (A) Total numbers of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations in the validation datasets (L (5min), maroon; H (5min), green) are shown as a bar plot. Bars for annotation-level totals for the human listener-labeled data are shown within the bar for the corresponding segment label, indicating that these are sub-categories. ChSp annotation types are C (canonical) and X (non-canonical non-reflexive), ChNsp annotation types are R (cry) and L (laugh), and Ad annotation types are T (infant-directed), U (addressee unknown), and N (not infant-directed). (B) Box plots summarizing total numbers of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations at the recording-level are shown for L (5min) and H (5min) data. (C) Box plots summarizing total vocalization numbers by annotation type (x-axis) at the recording-level are shown for H (5min) data. B and C share a y-axis. For H (5min) data, vocalization numbers are estimated after overlaps have been processed and for all datasets, number of vocalizations are reported after vocalizations with the same label are merged when separated by 0 s IEIs. For B and C, box plots whose notches do not overlap have different medians at the 0.05 significance level. Data plotted in B and C have not been normalized to account for differences in vocalization numbers at the recording-level due to variations in the number of annotated sections across recordings (see Section S2).
Figure S7: Vocalization durations in LENA daylong data by infant age and vocalization type, expressed as totals at the dataset-level and as box plots for totals at the recording day-level. (A) Total durations of infant (Ch; dark blue) and adult (Ad; green) vocalizations in the L (day) dataset for each infant age (months; x-axis) are shown as a bar plot. Bars for infant speech-related (ChSp) and infant non-speech-related (ChNsp) total durations are shown within the infant (Ch) bars, indicating that these are sub-categories. (B) Box plots summarizing total durations of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations at the recording day-level are shown as a function of infant age (months; x-axis). Box plots whose notches do not overlap have different medians at the 0.05 significance level. Note that total vocalization durations shown in A do not account for the fact that different ages have different numbers of daylong recordings (see Table S1). As such, B provides a more accurate representation of how the total duration of each vocalization type in a daylong recording changes with infant age. However, data plotted in B have not been normalized to account for differences in total vocalization durations at the recording-level due to variations in recording duration across daylong recordings (see Section S3.1).
Figure S8: Vocalization durations in the validation datasets by vocalization type, expressed as totals at the dataset-level and as box plots for totals at the recording-level. (A) Total durations of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations in the validation datasets (L (5min), maroon; H (5min), green) are shown as a bar plot. Bars for annotation-level totals for the human listener-labeled data are shown within the bar for the corresponding segment label, indicating that these are sub-categories. ChSp annotation types are C (canonical) and X (non-canonical non-reflexive), ChNsp annotation types are R (cry) and L (laugh), and Ad annotation types are T (infant-directed), U (addressee unknown), and N (not infant-directed). (B) Box plots summarizing total durations of infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations at the recording-level are shown for L (5min) and H (5min) data. (C) Box plots summarizing total vocalization durations by annotation type (x-axis) at the recording-level are shown for H (5min) data. B and C share a y-axis. For H (5min) data, vocalization durations are estimated after overlaps have been processed and for all datasets, vocalization durations are reported after vocalizations with the same label are merged when separated by 0 s IEIs. For B and C, box plots whose notches do not overlap have different medians at the 0.05 significance level. Data plotted in B and C have not been normalized to account for differences in total vocalization durations at the recording-level due to variations in the number of annotated sections across recordings (see Section S2).
S4 IEI and duration distributions of ChSp and Ad segments
Both IEI and duration distributions have long tails, with a large number of data points occurring in a relatively small range compared to the range of the data. We use a recursive binning approach based on percentile values (adapted from 3) to estimate (empirical) distributions for both IEI and duration data.
We first partition the data at a specified percentile value. For illustrative purposes, let us set this to the 5th percentile. Then, this process divides the data into two sets: the first consisting of all data less than or equal to the 5th percentile value (the ‘head’) and the second consisting of all data greater than the 5th percentile value (the ‘tail’). Since this is the first of several partitions, we refer to the head and tail obtained from this partition as the first head and the first tail. Next, we partition the first tail at its 5th percentile value to get the second tail. We repeat this process till partitioning is no longer possible—this is enforced by stopping partitioning when the resultant tail has fewer than 2 data points and discarding the corresponding partition value. Thus, we get a series of partitions for the data given by the set of 5th percentile values of the recursive tails. We use the partitions obtained using this method to bin the data and estimate the empirical distribution. The first bin ranges from the minimum value in the data to the 1st partition value (corresponding to the 5th percentile of the data), the second bin ranges from the 1st partition value to the 2nd partition value (corresponding to the 5th percentile of the first tail), and so on. The last bin ranges from the last partition value to the maximum value in the data. Finally, the bin centers are computed to plot the distributions.
To best represent the long-tailed nature of the data while also showcasing how the data is distributed in the tail, we present the data as a probability distribution as well as an empirical complementary cumulative distribution function (CCDF). The probability distribution is estimated by normalizing the bin counts such that the bin counts sum to 1. The CCDF is derived from the cumulative distribution function (CDF). The CDF estimates the probability that the value of a random variable, $X$, is less than or equal to a given number, $x$. This is expressed mathematically as $F_{X}(x)=P(X\leq x)$, where $F_{X}(x)$ is the CDF. The CCDF, then, is given by $1-F_{X}(x)=P(X>x)$, and provides an easy way to visualize distribution tails [4]. The empirical CCDFs presented in Sections S4.1 and S4.2 are derived from the probability distributions estimated using the method described above.
We use 5th percentile values to partition LENA daylong data, and 10th percentile values to partition validation data because of the much lower sample size of the latter. We also truncate the CCDFs such that probability values less than 10-6 are discarded.
IEI and duration distributions from L (day) and L (5min) data have artifacts resulting from minimum duration thresholds for different segment types as set by LENA™. Minimum duration thresholds for infant (speech-related and non-speech-related) and adult segments are 600 ms and 1 s, respectively. However, we find that roughly 8% of the adult vocalizations (after merging vocalizations of the same type separated by 0 s IEIs) in the L (day) data have durations ranging from 600 ms to 1 s. Accordingly, infant speech-related (ChSp) and adult (Ad) duration distributions from both L (day) and L (5min) data begin at or around 600 ms (see Section S4.2). Adult duration distributions peak at 1 ms—corresponding to the 1 s minimum—while ChSp duration distributions peak at 600 ms. Duration distributions for the human listener-labeled data, however, begin at lower duration values.
While LENA™ also sets minimum durations for all other segment labels (ranging from 800 ms to 1 s, depending on the segment type; see [2]), we note that the observed minimum duration for most segment types in the L (day) data—and by extension, L (5min) data—is 600 ms (Fig. S2B). As a result, the smallest possible IEIs for infant speech-related (ChSp) vocalizations (after merging vocalizations of the same type separated by 0 s IEIs) are necessarily 600 ms long—set by the minimum segment duration across all LENA segment labels—since at least one segment type other than ChSp must intervene between any two ChSp vocalizations. The same considerations also apply for adult IEIs. These lower limits are reflected in the IEI distributions for L (day) and L (5min) data but do not apply to human listener-labeled data (see Section S4.1).
S4.1 IEI distributions
Figure S9: Infant speech-related IEI distributions for LENA daylong data. (A) Empirical probability distributions of infant speech-related (ChSp) IEIs ($\Delta t$) for 3, 6, 9, and 18 months (see legend). (B) Empirical complementary cumulative distribution functions (CCDF) of ChSp IEIs for 3, 6, 9, and 18 months. Here, $P(\Delta T>\Delta t)$ indicates the probability that the random variable representing IEIs, $\Delta T$ is greater than the value of the IEI on the x-axis, $\Delta t$. A and B share a legend.
Figure S10: Infant speech-related IEI distributions for validation data. (A) Empirical probability distributions of infant speech-related (ChSp) IEIs ($\Delta t$) for H (5min; All Ad) and L (5min) data (see legend). (B) Empirical complementary cumulative distribution functions (CCDF) of ChSp IEIs for H (5min; All Ad) and L (5min) data. Here, $P(\Delta T>\Delta t)$ indicates the probability that the random variable representing IEIs, $\Delta T$ is greater than the value of the IEI on the x-axis, $\Delta t$. A and B share a legend.
Figure S11: Adult IEI distributions for LENA daylong data. (A) Empirical probability distributions of adult (Ad) IEIs ($\Delta t$) for 3, 6, 9, and 18 months (see legend). (B) Empirical CCDFs of Ad IEIs for 3, 6, 9, and 18 months. Here, $P(\Delta T>\Delta t)$ indicates the probability that the random variable representing IEIs, $\Delta T$ is greater than the value of the IEI on the x-axis, $\Delta t$. A and B share a legend.
Figure S12: Adult IEI distributions for validation data. (A) Empirical probability distributions of adult (Ad) IEIs ($\Delta t$) for H (5min; All Ad), H (5min; T-Ad), and L (5min) data (see legend). (B) Empirical CCDFs of Ad IEIs for H (5min; All Ad), H (5min; T-Ad), and L (5min) data. Here, $P(\Delta T>\Delta t)$ indicates the probability that the random variable representing IEIs, $\Delta T$ is greater than the value of the IEI on the x-axis, $\Delta t$. A and B share a legend.
S4.2 Duration distributions
Figure S13: Infant speech-related duration distributions for LENA daylong data. (A) Empirical probability distributions of infant speech-related (ChSp) durations ($d$) for 3, 6, 9, and 18 months (see legend). (B) Empirical CCDFs of ChSp durations for 3, 6, 9, and 18 months. Here, $P(D>d)$ indicates the probability that the random variable representing durations, $D$ is greater than the value of the duration on the x-axis, $d$. A and B share a legend.
Figure S14: Infant speech-related duration distributions for validation data. (A) Empirical probability distributions of infant speech-related (ChSp) durations ($d$) for H (5min; All Ad) and L (5min) data (see legend). (B) Empirical CCDFs of ChSp durations for H (5min; All Ad) and L (5min) data. Here, $P(D>d)$ indicates the probability that the random variable representing durations, $D$ is greater than the value of the duration on the x-axis, $d$. A and B share a legend.
Figure S15: Adult duration distributions for LENA daylong data. (A) Empirical probability distributions of adult (Ad) durations ($d$) for 3, 6, 9, and 18 months (see legend). (B) Empirical CCDFs of Ad durations for 3, 6, 9, and 18 months. Here, $P(D>d)$ indicates the probability that the random variable representing durations, $D$ is greater than the value of the duration on the x-axis, $d$. A and B share a legend.
Figure S16: Adult duration distributions for validation data. (A) Empirical probability distributions of adult (Ad) durations ($d$) for H (5min; All Ad), H (5min; T-Ad), and L (5min) data (see legend). (B) Empirical CCDFs of Ad durations for H (5min; All Ad), H (5min; T-Ad), and L (5min) data. Here, $P(D>d)$ indicates the probability that the random variable representing durations, $D$ is greater than the value of the duration on the X axis, $d$. A and B share a legend.
S5 Reliability between LENA and human listener labels
To get reliability estimates, we divided human and LENA-labeled 5-minute sections into 1 ms frames following [5] (see also Fig. S17). Human listener-labeled sections were compared against the corresponding LENA-labeled sections by lining up the start and end times of the 5-minute sections (Fig. S18). We used 1 ms frames as opposed to the choice of 10 ms frames in [5] to accommodate the resolution of human listener onset and offsets. We computed false alarm rates (FAR), miss rates (MR), confusion rates (CR), and identification error rates (IDER) as defined in [5] as well as simple percent agreement and Cohen’s kappa values using frame-by-frame comparisons between human listener-labels and LENA labels. We also computed precison and recall estimates as outlined in [5] and associated confusion matrices (Fig. S18). These reliability measures were computed for human listener-labeled data with all adult vocalizations included (H (5min; All Ad)), and human listener-labeled data with only child-directed adult vocalizations included (H (5min; T-Ad)). All reliability measures were computed at the day-level, i.e., aggregated over all 5-minute validation sections in a daylong recording.
Figure S17: Schematic describing labeling error rate estimation. A hypothetical sequence of human listener labels (indicated by ‘Hum’ to the left of the sequence) is shown to demonstrate processing steps prior to computing labeling error rates. Solid-tipped arrows indicate the sequence of processing steps. Human listeners identify onsets and offsets of key segments—infant speech-related (ChSp, light blue), infant non-speech-related (ChNsp, dark blue), and adult (Ad, gray)—and assign corresponding labels (see Methods for details). Note that the label sequence presented here is understood to have already gone through processing to remove overlaps and merge vocalizations of the same type when separated by 0 s IEIs (see Methods for details). Before estimating error rates, we assign ‘NA’ labels to all segments spanning times between labeled vocalizations (ChSp, ChNsp, Ad) as shown. While LENA provides exhaustive labels spanning the entire length of recorded audio, human listeners only code for key segments, and there is no information to compare against LENA labels that are not ChSp, ChNsp, or Ad. Therefore, for both LENA and human listener-labeled data, the NA label serves as a catch-all to indicate that the segment does not carry a key segment label. Segments labeled NA are indicated by the white spaces in the sequence. In the next step, the sequence is divided into ‘frames’ that are 1 ms long. The processing steps applied to human listener-labeled data are also applied to LENA-labeled data before estimating error rates.
Figure S18: Schematic describing labeling error rate estimation (continued). (A) The human listener-labeled sequence from Fig. S17 (indicated by ‘Hum’ to the left of the sequence) and the corresponding hypothetical LENA-labeled sequence (indicated by ‘LENA’ to the left of the sequence) are presented next to each other with 1 ms frames indicated. Frames with ChSp labels are light blue, frames with ChNsp labels are dark blue, frames with Ad labels are gray, and NA frames are white. The total number of frames identified by the human listener as containing key segments (T) is indicated above the human listener-labeled sequence. The number of frames with false alarms (FA), where LENA detected key segments but the human listener did not; misses (M), where LENA did not detect a key segment but the human listener did; and confusions (C), where LENA and the human listener detected key segments but disagreed on the key segment label, are indicated below the LENA sequence. (B) The confusion matrix for the human listener-labeled sequence and the corresponding LENA labels is shown (left side of panel B). The row indices, $i$, correspond to human listener labels (green, H) while column indices, $j$, correspond to LENA labels (red, L), both in order of ChSp, ChNsp, Ad, and NA. The $(i,j)$th element of the confusion matrix is the number of frames that were assigned the $i$th label by the human listener and the $j$th label by LENA. The diagonal, therefore, gives the number of frames that LENA and the human listener agreed on the label. The total number of frames labeled a certain type by the human listener is the sum of the corresponding row, while the total number of frames labeled a certain type by LENA is the sum of the corresponding column. The totals for each label type are shown as a separate column (human listener labels) and row (LENA labels) next to the confusion matrix. The overall total number of frames is the sum of the human total column or of the LENA total row, and is also indicated. The number of misses (M; orange), false alarms (FA; yellow), and confusions (C; gray), as well as the total number of frames identified by the human listener as key segments (T; purple outline) are indicated in or adjacent to the confusion matrix as applicable (see legend next to the matrix). On the left side of panel B, the miss rate (MR), false alarm rate (FAR), and confusion rate (CR) are computed for the label sequences presented in A. The identification error rate, which is the sum of MR, FAR, and CR, is also shown.
Figure S19: Confusion matrices for validation data. (A) Confusion matrix (expressed in seconds; see color bar) between LENA (x-axis) and human listener labels (y-axis) when all adult vocalizations are included in the human listener-labeled dataset, H (5min; All Ad). For example, the first row of the confusion matrix can be summarized as follows: LENA and human listeners agree on infant speech-related (ChSp) labels for 9659 s of audio, while 2900 s of audio labeled as ChSp by human listeners are labeled as infant non-speech-related (ChNsp) by LENA, 650 s of audio labeled as ChSp by human listeners are labeled as adult (Ad) by LENA, and 5676 s of audio labeled as ChSp by human listeners are identified as containing no key segment audio (corresponding to NA labels for reliability estimation purposes) by LENA. Note that the color bar is on a log scale. (B) Confusion matrix (expressed in seconds; see color bar) between LENA and human listener labels when only child-directed adult vocalizations are included in the human listener-labeled dataset, H (5min; T-Ad). As expected, only the last two rows in the confusion matrix (corresponding to human listener Ad and NA labels) differ from the confusion matrix in A. Further, the number of seconds of audio lost from the human listener-labeled Ad category (with respect to the confusion matrix in A) is gained by the NA category (also with respect to the confusion matrix in A). This is as expected, since only considering child-directed adult vocalizations results in an increase in seconds of audio that do not contain any key segment.
H (5min; All Ad) H (5min; T-Ad)
Reliability measure Mean Total Mean Total
Percent Agreement 72.61 72.61 72.94 72.99
Cohen’s Kappa 0.44 0.46 0.40 0.42
FAR 0.34 0.29 0.72 0.52
MR 0.36 0.35 0.32 0.31
CR 0.16 0.17 0.19 0.19
IER 0.87 0.81 1.23 1.02
Table S4: Reliability measures for validation data. Mean and total percent agreements, Cohen’s kappa values, false alarm rates (FAR), miss rates (MR), confusion rates (CR) and identification error rates (IDER) are shown when comparing human-listener data with all adult vocalizations included (H (5min, All Ad)) and only child-directed adult vocalizations included (H (5min; T-Ad)) against corresponding LENA-labeled 5 minute sections (L (5min)). The means are computed over recording day-level reliability measures where each file typically has up to three 5-minute sections (see Methods as well as Supplementary Section S2 for details).
Figure S20: Precision and Recall for key segments for validation data. (A) Total precision for key segments (infant speech-related, ChSp; infant non-speech-related, ChNsp; adult, Ad) between LENA and human listener labels are shown when all adult vocalizations are included in the human listener-labeled dataset (H (5min; All Ad); blue) vs. when only child-directed adult vocalizations are included, (H (5min; T-Ad); red). Precision is the number of frames LENA and human listeners agree on a label divided by the total number of frames LENA identifies as that label. High precision means that frames LENA identifies as a certain category are likely to have also been identified by human annotators as belonging to that category. (B) Total recall for key segments (ChSp, ChNsp, Ad) between LENA and human listener labels for H (5min; All Ad) and H (5min; T-Ad) data are shown. Recall is the number of frames LENA and human listeners agree on a label divided by the total number of frames human listeners identify as that label. High recall means that the frames human annotators identify as a category are likely to have also been labeled by LENA as that category. Precision and recall numbers for the two human listener-labeled datasets differ only for Ad vocalizations, since ChSp and ChNsp labels are the same for both human-labeled datasets.
S6 Methodological details
S6.1 Overlap processing for human listener-labeled data
Figure S21: Schematic describing overlap processing for human listener-labeled data. A hypothetical sequence of infant (blue) and adult (gray) vocalizations as labeled by human listeners are shown, prior to removing overlaps. Annotation labels for infant vocalizations (canonical, C; non-canonical non-reflexive, X; cry, R; or laugh, L) and adult vocalizations (infant-directed, T; not infant-directed, N; or addressee unknown, U) are indicated. The sequence of steps in the overlap processing protocol are indicated by the solid arrows. In the first step, overlaps (crosshatched portions) are identified based on onsets and offsets of vocalizations. In the second step, these overlaps are removed to obtain non-overlapping infant and adult vocalization time series. This results in vocalizations being shortened (e.g., first infant vocalization with annotation label C), removed (if they overlap fully with another vocalization; e.g., adult vocalization with annotation label U), or split into multiple vocalization events (e.g., adult vocalization with annotation label T). For vocalizations that are split into multiple vocalizations or in cases where part of a vocalization is removed as a result of overlap processing, the resultant non-overlapping or truncated vocalization(s) retain the segment and annotation labels of the ‘parent’ vocalization, as shown. Overlap processing was performed on the human listener-labeled dataset with all adult vocalizations included (H (5min; All Ad)) prior to extracting the human listener-labeled dataset with only child-directed adult vocalizations included (H (5min; T-Ad)).
ChSp ChNsp Ad
Number of full overlaps 1471 124 534
Number of partial overlaps 2666 562 3668
Total vocalization number before overlap removal 25522 2600 17458
Total vocalization number after overlap removal 24341 2665 18437
Total vocalization duration (s) before overlap removal 21050.18 4355.82 22937.70
Total vocalization duration (s) after overlap removal 18928.32 3670.82 20130.84
Table S5: A summary of overlap processing: vocalization numbers and durations. The number of full and partial overlaps as well as total numbers and durations of vocalizations before and after removing overlaps are shown for infant speech-related (ChSp), infant non-speech-related (ChNsp), and adult (Ad) vocalizations. A full overlap is when the entire vocalization in question is removed as part of overlap processing (e.g., adult vocalization with annotation label U in Fig. S21). A partial overlap is when part of the vocalization in question is removed during overlap processing (e.g., both infant vocalizations with annotation label R in Fig. S21).
S6.2 Merging vocalizations separated by 0 s IEIs
It was possible for both LENA™ and human listener labels to have vocalizations of the same type perfectly adjacent to each other, with no time between the offset of one vocalization and the onset of the next vocalization of the same type. These were identifiable in our analysis pipeline as cases of 0 s IEIs. Prior to all statistical analyses (and in the case of human listener-labeled data, after overlap processing), we merged all vocalizations of the same type separated by 0 s IEIs. While most of these merges involved a pair of vocalizations being merged into one vocalization, there were also cases where a sequence of more than two vocalizations separated by 0 s IEIs was merged into one vocalization. We refer to the latter as a merge ‘chain’.
ChSp ChNsp Ad
L (day) 10950 5897 47197
L (5min) 1083 257 816
H (5min: All Ad) 8 5 141
H (5min; T-Ad) 8 5 15
Table S6: A summary of 0 s IEI merges. The number of instances where 2 or more vocalizations of the same type separated by 0 s IEI(s) were merged into one vocalization are shown by vocalization type (infant speech-related, ChSp; infant non-speech-related, ChNsp; adult, Ad) for LENA daylong data (L (day)) as well as validation data (human listener-labeled data with all adult vocalizations included, H (5min; All Ad); human listener-labeled data with only infant-directed adult vocalizations included, H (5min; T-Ad); and corresponding LENA-labeled 5-minute sections, L (5min)). For a summary of the number of such vocalization merges involving more than 2 vocalizations, see below.
Figure S22: Frequency distributions of vocalization merge chains. (A) The number of instances (y-axis) of vocalization merge chains—where more than 2 vocalizations of the same type separated by 0 s IEIs were merged into one vocalization—are shown as a function of the number of vocalizations in a merge chain (x-axis), for infant speech-related (ChSp, blue), infant non-speech-related (ChNsp, red), and adult (Ad, yellow) vocalizations in LENA daylong (L (day)) data. (B) Number of vocalization merge chains as a function of the number of vocalizations in a merge chain are shown for ChSp, ChNsp, and Ad vocalizations in the LENA-labeled validation subset (L (5min) data). A and B share a legend.
Unlike both LENA-labeled datasets (L (day) and L (5min)), all vocalization merge chains for human listener-labeled data had three vocalizations. For H (5min; All Ad) data, there were 10 such merge chains for adult vocalizations, while for H (5min; T-Ad) data, there was only one merge chain for adult vocalizations. For both H (5min; All Ad) and H (5min; T-Ad) datasets, there was one vocalization merge chain for ChSp vocalizations and one merge chain for ChNsp vocalizations.
S6.3 Response computation and analysis
Figure S23: A demonstration of the effect of IEI transformation on data distribution using infant IEIs at 18 months. (A) A scatter plot of current infant speech-related (ChSp) IEIs (IEIi) plotted against previous ChSp IEIs (IEIi-1) is shown using LENA daylong data (L (day)) from infants at 18 months. The data presented here is the same as that in Fig. 1 in the main text; however, IEIs depicted here are ‘raw’ IEIs, prior to log-transformation and normalization. To remain consistent with Fig. 1 in the main text, the plot displays raw IEIs corresponding to the same 2000 data points (chosen randomly from the full set of daylong LENA infant IEIs at 18 months) in Fig. 1C and D (see main text). (B) A scatter plot for the same data after IEIs were log-transformed and normalized with respect to the full 18 month L (day) infant IEI dataset is shown. This is a reproduction of Fig. 1C in the main text. The normalization is indicated by the $f_{z}$ function. The dashed line represents the regression line obtained by regressing (log-transformed and normalized) current IEIs over previous IEIs within the 18 month daylong infant data, and is used to compute current IEI residuals which go into our response analyses (see Methods).
Figure S24: A demonstration of the effect of residual estimation on data distribution for Resp and NoResp IEIs using infant IEIs at 18 months. (A) shows infant speech-related (ChSp) IEI distributions at 18 months for LENA daylong data, prior to log-transformation and normalization. Data has been separated into IEIs associated with an adult response (Resp; green) and IEIs not associated with an adult response (NoResp; red) for a response window ($T_{\rm{resp}}$) value of 5 s. IEIs associated with NA responses are not represented. (B) uses the same data to show the distributions of Resp (green) and NoResp (red) IEIs after log-transformation and normalization. (C) uses the same data to show the distributions of Resp (green) and NoResp (red) IEI residuals. In A, B, and C, each distribution has been normalized independently.
Figure S25: Infant Resp and NoResp IEI residual distributions for LENA daylong data for $T_{\rm resp}$ = 5s. (A) Distributions of infant speech-related (ChSp) IEI residuals separated into IEIs associated with an adult response (Resp; green) and IEIs not associated with an adult response (NoResp; green) are shown for LENA daylong data using data from infants at 3 months, for $T_{\rm resp}$ = 5s. Residuals for IEIs associated with NA responses are not shown. Resp and NoResp distributions have been normalized independently. (B), (C), and (D) show similar plots for Resp and NoResp ChSp IEI residuals for infants at 6, 9, and 18 months, respectively. All plots share a legend. The number of bins for each distribution was determined separately using the Freedman-Diaconis rule [6]. As such, not all distributions have the same number of bins.
Figure S26: Adult Resp and NoResp IEI residual distributions for LENA daylong data for $T_{\rm resp}$ = 5s. (A) Distributions of adult IEI residuals separated into IEIs associated with an infant response (Resp; green) and IEIs not associated with an infant response (NoResp; green) are shown for LENA daylong data using data when infants are 3 months old, for $T_{\rm resp}$ = 5s. Residuals for IEIs associated with NA responses are not shown. Resp and NoResp distributions have been normalized independently. (B), (C), and (D) show similar plots for Resp and NoResp adult IEI residuals when infants are 6, 9, and 18 months old, respectively. All plots share a legend. The number of bins for each distribution was determined separately using the Freedman-Diaconis rule [6]. As such, not all distributions have the same number of bins.
Figure S27: Number of IEIs associated with different response types for LENA daylong data. (A) Total number of infant speech-related (ChSp) and adult (Ad) IEIs are shown as a bar plot. (B) Fractions of ChSp IEIs associated with an adult response (blue, Y response), not associated with an adult response (red, N response), and associated with NA response (yellow) are shown as a function of response window, $T_{\rm{resp}}$ (x-axis, seconds). All fractions are computed with respect to the total number of ChSp IEIs. (C) Fractions of Ad IEIs associated with a ChSp response (blue, Y response), not associated with a ChSp response (red, N response), and associated with NA response (yellow) are shown as a function of response window, $T_{\rm{resp}}$. All fractions are computed with respect to the total number of Ad IEIs. Note that B and C share the y-axis and legend.
Figure S28: Total number of validation data IEIs by vocalization type. The total number of infant speech-related (ChSp) and adult (Ad) IEIs are shown as bar plots for the validation data: human listener-labeled 5-minute sections with all adult vocalizations (H (5min; All-Ad); blue); human listener-labeled 5-minute sections with only child-directed adult vocalizations included (H (5min; T-Ad); green); and the corresponding LENA subset of 5-minute sections (L (5min); yellow). Note that the number of ChSp IEIs are the same for H (5min; All Ad) and H (5min; T-Ad) data since only the adult vocalizations differ between those datasets.
Figure S29: Number of IEIs associated with different response types for validation data. (A) Fractions of infant speech-related (ChSp) IEIs associated with an adult (Ad) response (blue, Y response), not associated with an adult response (red, N response), and associated with NA response (yellow) are shown as a function of response window, $T_{\rm{resp}}$ (x-axis, seconds), for L (5min) (solid line, open hexagrams), H (5min; All Ad) (dotted line, open diamonds), and H (5min; T-Ad) (dot-dash line, solid circles) data. All fractions are computed with respect to the total number of ChSp IEIs for that data type. For example, the fraction of L (5min) ChSp IEIs associated with a Y response is computed with respect to the total number of ChSp IEIs for L (5min) data. (B) Fractions of Ad IEIs associated with a ChSp response (blue, Y response), not associated with a ChSp response (red, N response), and associated with NA response (yellow) are shown as a function of response window, $T_{\rm{resp}}$, for L (5min) (solid line, open hexagrams), H (5min; All Ad) (dotted line, open diamonds), and H (5min; T-Ad) (dot-dash line, solid circles) data. All fractions are computed with respect to the total number of Ad IEIs for that data type. A and B share the y-axis and legend.
S7 Extended results
S7.1 Correlations between current and previous IEI
Infant age ChSp previous IEI $\beta$ Ad previous IEI $\beta$
3 months 0.30 0.33
6 months 0.26 0.31
9 months 0.25 0.31
18 months 0.23 0.29
Table S7: Previous IEI $\beta$s for LENA daylong data. Standardized regression coefficients ($\beta$s) from linear mixed effects analyses testing the effect of previous IEI on current IEI, with infant ID as a random effect are shown for LENA daylong data, by vocalization type and infant age. All effects are significant at the $p<0.001$ level.
Data type ChSp previous IEI $\beta$ Ad previous IEI $\beta$
L (5min) 0.10 0.17
H (5min; All Ad) 0.19 0.17
H (5min; T-Ad) 0.19 0.12
Table S8: Previous IEI $\beta$s for validation data. Standardized regression coefficients ($\beta$s) from linear mixed effects analyses testing the effect of previous IEI on current IEI, with infant ID, infant age, and infant ID-infant age interaction as random effects are shown for validation datasets, by vocalization type. All effects are significant at the $p<0.001$ level.
Data type Voc type Age $\beta$ Age2 $\beta$
L (day) ChSp -0.013∗∗∗ 4 × $\times\>10^{-4}$
($p<$ 0.001) ($p$ = 0.01)
L (5min) ChSp 0.006 -2 × $\times\>10^{-4}$
($p$ = 0.54) ($p$ = 0.52)
H (5min; All Ad) ChSp -0.014 6 × $\times\>10^{-4}$
($p$ = 0.14) ($p$ = 0.17)
H (5min; T-Ad) ChSp -0.014 6 × $\times\>10^{-4}$
($p$ = 0.14) ($p$ = 0.17)
L (day) Ad -0.003 5 × $\times\>10^{-5}$
($p$ = 0.31) ($p$ = 0.73)
L (5min) Ad -0.009 4 × $\times\>10^{-4}$
($p$ = 0.56) ($p$ = 0.63)
H (5min; All Ad) Ad -0.005 8 × $\times\>10^{-5}$
($p$ = 0.65) ($p$ = 0.87)
H (5min; T-Ad) Ad 0.004 -4 × $\times\>10^{-4}$
($p$ = 0.85) ($p$ = 0.67)
Table S9: Effect of infant age on previous IEI $\beta$s. Standardized regression coefficients ($\beta$s) from linear mixed effects analyses testing the effect of infant age on previous IEI $\beta$s are shown for infant speech-related (ChSp) and adult (Ad) vocalizations for LENA daylong data (L (day)) and validation data: human listener-labeled 5-minute sections with all adult vocalizations (H (5min; All-Ad)); human listener-labeled 5-minute sections with only child-directed adult vocalizations included (H (5min; T-Ad)); and the corresponding LENA subset of 5-minute sections (L (5min)). Linear mixed effects analyses were run predicting recording-level previous IEI $\beta$s—correlations between current IEI and previous IEI at the recording level—with infant age and infant age2 as fixed effects and infant ID as a random effect. Significant $\beta$s at $p=0.05$ are in bold while significant $\beta$s at $p=0.001$ are indicated by ***.
S7.2 Response effect on IEIs
Tables with response effect $\beta$s using a response window ($T_{\rm resp}$) value of 5 s as a representative case are shown below. Response $\beta$s for the effect of adult (Ad) responses on infant speech-related (ChSp) IEIs and for the effect of infant speech-related responses on adult IEIs are shown for LENA daylong data and validation data. Results for the full range of $T_{\rm{resp}}$ values analyzed are available in the OSF repository associated with this study.
Infant age ChSp IEI Response $\beta$ 99.9% CI
3 months -0.12 (-0.16, -0.08)
6 months -0.09 (-0.13, -0.05)
9 months -0.11 (-0.15, -0.07)
18 months -0.15 (-0.19, -0.11)
Table S10: Effect of receiving an adult response on infant speech-related IEI length for LENA daylong data for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving an Ad response on ChSp IEI residuals (after regressing current IEI on previous IEI with infant ID as a random effect) for $T_{\rm resp}=5$ s are shown, by infant age. 99.9% confidence intervals for the $\beta$s are also provided. All effects are significant at the $p<0.001$ level.
Infant Age Ad IEI Response $\beta$ 99.9% CI
3 months -0.17 (-0.21, -0.12)
6 months -0.11 (-0.15, -0.06)
9 months -0.14 (-0.19, -0.10)
18 months -0.13 (-0.17, -0.09)
Table S11: Effect of receiving an infant response on adult IEI length for LENA daylong data for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving a ChSp response on Ad IEI residuals (after regressing current IEI on previous IEI with infant ID as a random effect) for $T_{\rm resp}=5$ s are shown, by infant age. 99.9% confidence intervals for the $\beta$s are also provided. All effects are significant at the $p<0.001$ level.
Data type ChSp IEI Response $\beta$ 99.9% CI
L (5min) -0.07 ($p$ = 0.05) (-0.18, 0.05)
H (5min; All Ad) -0.03 ($p$ = 0.30) (-0.14, 0.07)
H (5min; T-Ad) -0.07 ($p$ = 0.03) (-0.19, 0.04)
Table S12: Effect of receiving an adult response on infant speech-related IEI length for validation datasets for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving an Ad response on ChSp IEI residuals (after regressing current IEI on previous IEI with infant ID, infant age, and infant ID-infant age interaction as random effects) for $T_{\rm resp}=5$ s are shown for the validation datasets: human listener-labeled 5-minute sections with all adult vocalizations (H (5min; All-Ad)); human listener-labeled 5-minute sections with only child-directed adult vocalizations included (H (5min; T-Ad)); and the corresponding LENA subset of 5-minute sections (L (5min)). $p$ values are indicated in parentheses. Significant $\beta$s at $p$ = 0.05 are in bold (the $p$ value for L (5min) data is 0.0478 and has been rounded to 0.05). 99.9% confidence intervals for the $\beta$s are also provided.
Data type Ad IEI Response $\beta$ 99.9% CI
L (5min) -0.12 ($p$ = 0.002)** (-0.25, 0.01)
H (5min; All Ad) 0.04 ($p$ = 0.30) (-0.09, 0.17)
H (5min; T-Ad) -0.16*** ($p<$ 0.001) (-0.31, -0.001)
Table S13: Effect of receiving an infant response on adult IEI length for validation datasets for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving a ChSp response on Ad IEI residuals (after regressing current IEI on previous IEI with infant ID, infant age, and infant ID-infant age interaction as random effects) for $T_{\rm resp}=5$ s are shown for the validation datasets. $p$ values are indicated in parentheses. Significant $\beta$s at $p$ = 0.05 are in bold, significant $\beta$s at $p$ = 0.01 are indicated by **, and significant $\beta$s at $p$ = 0.001 are indicated by ***. 99.9% confidence intervals for the $\beta$s are also provided.
For tables summarizing the effect of infant age on recording level response $\beta$s (none significant at $p<0.001$), see the OSF repository associated with this project.
[…]